It says perform the indicated operation. How do I do this
(x+2)(x^2-3x+8)
Can someone help?

Answer by queenofit(120)

I do these different than most people. You are multiplying these numbers together, so use x+2 and multiply the x by each number/variable in the second equation then multiply the 2 by each number/variable in the second equation. then do the adding and subtraction of the common factors.
like I said I do it different, I do it like long multiplication, it makes it easier for me and I will show you how...maybe it will help...
x^2-3x+8 multiply just like you would
* x+2 with all numbers in the equation
__________
2x^2-6x+16 make sure you line up with the
x^3 -3x^2+8x same type of number or variable
+__________________
x^3-x^2+2x+16
you can do this problem in the other form that is usually taught,
(x+2)(x^2-3x+8)
take the x and multiply each number in the second equation by x
x*x^2 is x^3 x*(-3x)=-3x^2 x*8=8x
by 2
x^2*2=2x^2 2*(-3x)=-6x x*8=16
so you have x^3-3x^2-8x+2x^2-6x+16
combine the common factors and you have x^3-3x^2+2x^2-6x+8x+16=
x^3-x^2+2x+16